A radiator contains 7 qt of a 30% antifreeze solution. How much should be removed and replaced with pure antifreeze to yield a 40% solution?
Answer by amoresroy(361)

Let x = the amount of solution to be replaced with pure antifreeze
0.3(7-x) + x = .4(7)
2.1 -.3x + x = 2.8
.7x = 2.8-2.1
x = 1 qt